# FG-CLIP 2: A Bilingual Fine-grained Vision-Language Alignment Model

### v2

@@ 4.2.1 Fine-Grained Understanding
We evaluate open-source image-text alignment models on FG-OVD (Bianchi et al., 2024), a fine-grained benchmark emphasizing grounding in specific local image regions. Each region is paired with one positive description and ten synthetically perturbed negatives, forming a challenging discrimination task. The benchmark comprises four subsets: trivial, easy, medium, and hard, arranged by increasing linguistic subtlety between correct and distractor texts, requiring finer-grained reasoning for accurate matching. Following FineCLIP (Jing et al., 2024), we extract dense visual features and use ROIAlign with provided bounding boxes to obtain region-specific representations. Similarity scores are computed between region features and textual descriptions, with Top-1 accuracy used as the evaluation metric. Results on the left side of Table 1 show that FG-CLIP 2 achieves significant gains over prior models. This demonstrates its superior capability in distinguishing subtle visual-linguistic correspondences, a key requirement for fine-grained understanding.
@@
4.2.2 Bounding Box Classification
We evaluate zero-shot bounding box classification on COCO-val2017 (Lin et al., 2014), LVIS (Gupta et al., 2019), and our proposed BoxClass-CN dataset, following the protocol of (Xie et al., 2025). While the former focus on English category recognition within localized regions, BoxClass-CN targets Chinese, enabling a bilingual assessment of fine-grained vision-language alignment. As shown in the right part of Table 1, FG-CLIP 2 achieves state-of-the-art performance in both languages and significantly outperforms all compared open-source models. These results demonstrate its robust ability to align local visual content with semantic concepts across linguistic boundaries.
@@ 4.2.3 Open-Vocabulary Object Detection
To assess the impact of vision-language alignment models on open-vocabulary detection (OVD), we adopt a training-free evaluation protocol that avoids biases from detector fine-tuning. Unlike prior approaches (Wu et al., 2024) that require small-scale detector training on fixed categories, our method directly leverages the zero-shot generalization of alignment models to calibrate confidence scores and category predictions in the final OVD output. This allows for a cleaner analysis of the contribution of image-text alignment to OVD performance. We adopt a fusion strategy that leverages the vision-language alignment model to refine the class predictions and confidence scores of a pre-trained detector. Specifically, similarity scores from the alignment model are combined with the detector’s original confidences via geometric averaging, enabling more semantically accurate and calibrated open-vocabulary detection. We use LLMDet (Fu et al., 2025) as the base detector and evaluate on the challenging LVIS dataset (Gupta et al., 2019), which contains 1,203 categories. Categories are split into frequent, common, and rare based on occurrence frequency, and AP is reported per group. As shown in Table 2, FG-CLIP 2 combined with LLMDet achieves the best performance among open-source methods, demonstrating strong practical utility and validating its superior generalization in detection scenarios.
@@ 4.3.1 Long/Short Caption Image-Text Retrieval
To comprehensively evaluate image-text alignment under varying linguistic complexity, we conduct experiments on both short and long caption retrieval tasks. Short-text retrieval (Flickr30k (Young et al., 2014), MSCOCO (Lin et al., 2014)) assesses basic semantic matching, while long-text retrieval requires fine-grained understanding of detailed descriptions and complex visual scenes. For long-text retrieval in English, we use the 1K subset of ShareGPT4V (Chen et al., 2024b) and the full test set of DCI (Urbanek et al., 2024) following the protocol of Long-CLIP (Zhang et al., 2024a). For Chinese, we evaluate on LIT-CN, DCI-CN, and DOCCI-CN for long-text, and Flickr-CNA (Xie et al., 2023) and COCO-CN (Li et al., 2019) for short-text. These datasets cover diverse content and caption styles, providing a robust evaluation of multilingual performance. As shown in Table 3 and Table 4, FG-CLIP 2 achieves consistent improvements across all settings, with particularly strong gains on long-text retrieval, highlighting its superior capability in fine-grained vision-language alignment. Notably, FG-CLIP 2 outperforms Meta CLIP 2 (Chuang et al., 2025), the current multilingual SOTA, on both language settings, despite using a smaller ViT-L/16 backbone with 1.0 billion parameters compared to Meta CLIP 2’s 1.8 billion parameter ViT-H/14. This highlights the effectiveness of our training paradigm in achieving stronger performance with reduced model scale.
4.3.2 Zero-shot Image Classification
We evaluate zero-shot image classification on ImageNet-1K (Deng et al., 2009) and ImageNet-v2 (Recht et al., 2019) using standard prompts (Radford et al., 2021). As shown in Table 3, FG-CLIP 2 achieves competitive performance compared to SigLIP 2 (Tschannen et al., 2025), and outperforms EVA-CLIP (Sun et al., 2023), Long-CLIP (Zhang et al., 2024a), and FineCLIP (Jing et al., 2024). This confirms that the improvements in fine-grained alignment do not come at the cost of standard recognition accuracy, demonstrating a well-balanced representation capability.
@@ 4.4 Dense prediction tasks
Table 6: Comparisons on large multimodal model benchmarks.
Method GQA MMMU TextVQA RefCOCO MMBench-EN MMBench-CN
val testA testB dev test dev test
LLaVA-1.5 + CLIP 61.9 35.7 58.2 76.2 83.4 67.9 65.1 66.5 58.2 58.4
LLaVA-1.5 + SigLIP 2 62.0 37.0 56.6 79.5 84.4 73.9 66.2 64.8 58.0 57.2
LLaVA-1.5 + Meta CLIP 2 62.4 37.0 59.4 76.7 82.8 69.8 66.5 67.0 59.3 60.3
LLaVA-1.5 + FG-CLIP 2 64.0 38.1 62.0 84.9 89.8 79.5 67.6 66.5 60.5 61.4
Table 7: Ablation study results for FG-CLIP 2.
Method Flickr30k ImageNet-V2 COCO80 FG-OVD
I2T T2I Top-1 Top-1 Top-5 Hard Medium Easy Trivial
FG-CLIP 2 94.1 81.9 72.2 74.9 95.7 52.3 76.3 80.3 92.0
w/o $\mathcal{L}_{\text{CMR}}$ 93.3 81.9 72.1 74.0 94.9 50.9 75.1 77.8 93.5
w/o $\mathcal{L}_{\text{TIC}}$ 93.7 81.8 72.1 70.1 94.8 51.6 75.1 79.1 92.1
w/o $\mathcal{L}_{\text{CMR}}$, $\mathcal{L}_{\text{TIC}}$ 93.5 81.6 72.0 62.7 93.4 51.7 76.0 77.6 90.7
@@ 4.5 Large Multimodal Model tasks
We empoly FG-CLIP 2 as a vision encoder in large multimodal models (LMMs) to assess its compatibility and utility in advanced multimodal reasoning. We integrate FG-CLIP 2 into a standard LLaVA-style LMM architecture, following the pre-training and supervised fine-tuning protocol of LLaVA-1.5 (Liu et al., 2023a). Results in Table 6 show that LMMs equipped with FG-CLIP 2 outperform those using other open-source encoders, including SigLIP 2 and Meta CLIP 2. This indicates that the fine-grained and bilingual capabilities of FG-CLIP 2 effectively transfer to higher-level multimodal tasks, making it a strong candidate for integration into next-generation LMMs.
@@ 4.6 Ablation Study
As our training framework leverages global-level, region-level, and hard-negative data through complementary learning mechanisms, we adopt a baseline model that incorporates global alignment alongside fine-grained visual and textual learning. This allows us to specifically evaluate the impact of the Cross-modal Rank Loss with Global Threshold Synchronization ($\mathcal{L}_{\text{CMR}}$) and the Textual Intra-modal Contrastive Loss ($\mathcal{L}_{\text{TIC}}$). As shown in Table 7, removing $\mathcal{L}_{\text{TIC}}$ leads to a 4.8-point drop in COCO Top-1 accuracy (to 70.1%) and a decrease in FG-OVD Hard performance from 52.3% to 51.6%, confirming its critical role in distinguishing semantically similar texts. Removing $\mathcal{L}_{\text{CMR}}$ also degrades performance, with FG-OVD Hard falling to 50.9%, indicating its importance in cross-modal alignment. When both losses are removed, COCO Top-1 drops sharply to 62.7%, demonstrating their complementary benefits. The full model achieves consistent gains across all tasks, validating the effectiveness of our proposed training objectives.
@@ 5 Conclusion and Limitations
In this work, we present FG-CLIP 2, a bilingual vision-language model that advances fine-grained understanding for both English and Chinese. Our two-stage training paradigm progressively refines alignment by leveraging both short and long captions, region-text supervision, and multiple discriminative objectives, including the proposed Textual Intra-modal Contrastive (TIC) loss to better distinguish semantically similar descriptions. Trained on large-scale, high-quality English and Chinese datasets, FG-CLIP 2 achieves superior performance across 29 datasets and 8 tasks, demonstrating strong bilingual generalization. To advance evaluation in non-English settings, we introduce a new benchmark for Chinese multimodal understanding with challenging tasks such as long caption image-text retrieval and bounding box classification. We release the model, code, and benchmark to support future research in bilingual fine-grained vision-language understanding. In future work, we will focus on extending the model to handle longer textual inputs and explicitly model relational structures among objects, enabling richer fine-grained multimodal understanding.
@@ Appendix A Training Data Details
Table A summarizes the datasets used in our two-stage training pipeline.
@@
In-house Data Chinese 12M
Appendix B Open-Vocabulary Object Detection Experimental Details
For each detected bounding box, we extract its visual feature by applying RoI-Align on the dense ViT feature map produced by FG-CLIP 2. We also encode all candidate category names into text embeddings using the same model. We then compute the cosine similarity between the region’s visual feature and each category’s text embedding, followed by Softmax normalization to obtain a category-wise alignment similarity distribution. To produce the final prediction, we combine LLMDet’s original confidence score with FG-CLIP 2’s normalized similarity score through a weighted multiplicative fusion. The resulting fused score reflects both the detector’s localization confidence and the alignment model’s semantic relevance. We then select the category with the highest fused score as the final predicted class, and assign the corresponding fused value as the final confidence output.

### v3

@@ 4.1 Implementation Details
Table 1: Performance comparison on fine-grained understanding tasks using Top-1 accuracy.
Method Backbone Fine-Grained Understanding
Hard Medium Easy Trivial
CLIP ViT-B/16 12.0 23.1 22.2 58.5
EVA-CLIP ViT-B/16 14.0 30.1 29.4 58.3
Long-CLIP ViT-B/16 9.2 18.4 16.2 51.8
FineCLIP ViT-B/16 26.8 49.8 50.4 71.9
SigLIP 2 ViT-B/16 24.9 46.5 48.7 85.0
FG-CLIP ViT-B/16 46.1 66.6 68.7 83.4
FG-CLIP 2 ViT-B/16 52.3 76.3 80.3 92.0
CLIP ViT-L/14 15.4 25.3 25.7 38.8
EVA-CLIP ViT-L/14 18.3 38.4 35.2 62.7
Long-CLIP ViT-L/14 9.6 19.7 16.0 39.8
FineCLIP ViT-L/14 22.8 46.0 46.0 73.6
SigLIP 2 ViT-L/16 24.1 47.1 47.4 84.1
FG-CLIP ViT-L/14 48.4 69.5 71.2 89.7
FG-CLIP 2 ViT-L/16 55.5 77.5 83.1 92.5
Meta CLIP 2 ViT-H/14 16.5 36.6 34.7 79.6
SigLIP 2 ViT-So/16 26.0 48.7 49.9 87.4
FG-CLIP 2 ViT-So/16 54.0 77.4 79.8 93.5
Table 2: Performance comparison on bounding box classification tasks using Top-1 accuracy.
Method Backbone COCO80 LVIS1203 BoxClass-CN566
CLIP ViT-B/16 44.2 20.9 -
EVA-CLIP ViT-B/16 30.6 14.4 -
Long-CLIP ViT-B/16 36.7 18.2 -
FineCLIP ViT-B/16 48.4 23.3 -
SigLIP 2 ViT-B/16 53.4 20.6 57.9
FG-CLIP ViT-B/16 52.3 28.6 -
FG-CLIP 2 ViT-B/16 74.9 47.3 60.7
CLIP ViT-L/14 33.8 9.3 -
EVA-CLIP ViT-L/14 32.1 18.3 -
Long-CLIP ViT-L/14 35.6 10.4 -
FineCLIP ViT-L/14 54.5 22.5 -
SigLIP 2 ViT-L/16 54.7 25.9 56.6
CLOC ViT-L/14 72.9 32.6 -
FG-CLIP ViT-L/14 63.2 38.3 -
FG-CLIP 2 ViT-L/16 74.0 41.9 68.6
Meta CLIP 2 ViT-H/14 52.0 24.4 55.2
SigLIP 2 ViT-So/16 62.0 31.4 63.6
FG-CLIP 2 ViT-So/16 77.4 43.9 66.5
@@ 4.2.1 Fine-Grained Understanding
We evaluate open-source image-text alignment models on FG-OVD (Bianchi et al., 2024), a fine-grained benchmark emphasizing grounding in specific local image regions. Each region is paired with one positive description and ten synthetically perturbed negatives, forming a challenging discrimination task. The benchmark comprises four subsets: trivial, easy, medium, and hard, arranged by increasing linguistic subtlety between correct and distractor texts, requiring finer-grained reasoning for accurate matching. Following FineCLIP (Jing et al., 2024), we extract dense visual features and use ROIAlign with provided bounding boxes to obtain region-specific representations. Similarity scores are computed between region features and textual descriptions, with Top-1 accuracy used as the evaluation metric. Results in Table 1 show that FG-CLIP 2 achieves significant gains over prior models. This demonstrates its superior capability in distinguishing subtle visual-linguistic correspondences, a key requirement for fine-grained understanding.
@@
4.2.2 Bounding Box Classification
We evaluate zero-shot bounding box classification on COCO-val2017 (Lin et al., 2014), LVIS (Gupta et al., 2019), and our proposed BoxClass-CN dataset, following the protocol of (Xie et al., 2025). While COCO and LVIS focus on English category recognition within localized regions, BoxClass-CN targets Chinese, enabling a bilingual assessment of fine-grained vision-language alignment. As shown in Table 2, FG-CLIP 2 achieves state-of-the-art performance in both languages and significantly outperforms all compared open-source models. These results demonstrate its robust ability to align local visual content with semantic concepts across linguistic boundaries.
@@ 4.2.3 Open-Vocabulary Object Detection
To assess the impact of vision-language alignment models on open-vocabulary detection (OVD), we adopt a training-free evaluation protocol that avoids biases from detector fine-tuning. Unlike prior approaches (Wu et al., 2024) that require small-scale detector training on fixed categories, our method directly leverages the zero-shot generalization of alignment models to calibrate confidence scores and category predictions in the final OVD output. This allows for a cleaner analysis of the contribution of image-text alignment to OVD performance. We adopt a fusion strategy that leverages the vision-language alignment model to refine the class predictions and confidence scores of a pre-trained detector. Specifically, similarity scores from the alignment model are combined with the detector’s original confidences via geometric averaging, enabling more semantically accurate and calibrated open-vocabulary detection. We use LLMDet (Fu et al., 2025) as the base detector and evaluate on the challenging LVIS dataset (Gupta et al., 2019), which contains 1,203 categories. As shown in Table 3, FG-CLIP 2 combined with LLMDet achieves the best performance among open-source methods, demonstrating its strong practical utility and superior generalization in detection scenarios.
@@ 4.3.1 Long/Short Caption Image-Text Retrieval
To comprehensively evaluate image-text alignment under varying linguistic complexity, we conduct experiments on both short and long caption retrieval tasks. Short-text retrieval (Flickr30k (Young et al., 2014), MSCOCO (Lin et al., 2014)) assesses basic semantic matching, while long-text retrieval requires fine-grained understanding of detailed descriptions and complex visual scenes. For short-text retrieval, we employ the validation set of MSCOCO and the test set of Flickr30k, which are widely used benchmarks for assessing image-text alignment models. For long-text retrieval in English, we use the 1K subset of ShareGPT4V (Chen et al., 2024b) and the full test set of DCI (Urbanek et al., 2024) following the protocol of Long-CLIP (Zhang et al., 2024a). For Chinese, we evaluate on LIT-CN, DCI-CN, and DOCCI-CN for long-text, and Flickr-CNA (Xie et al., 2023) and COCO-CN (Li et al., 2019) for short-text. We employ the validation set of COCO-CN and the test set of Flickr-CNA. These datasets cover diverse content and caption styles, providing a robust evaluation of multilingual performance. As shown in Table 4 and Table 5, FG-CLIP 2 achieves consistent improvements across all settings, with particularly strong gains on long-text retrieval, highlighting its superior capability in fine-grained vision-language alignment. Notably, FG-CLIP 2 outperforms Meta CLIP 2 (Chuang et al., 2025), the current multilingual SOTA, on both language settings, despite using a smaller ViT-L/16 backbone with 1.0 billion parameters compared to Meta CLIP 2’s 1.8 billion parameter ViT-H/14. This highlights the effectiveness of our training paradigm in achieving stronger performance with reduced model scale.
@@
Table 7: Comparisons on large multimodal model benchmarks.
Method GQA MMMU TextVQA RefCOCO MMBench-EN MMBench-CN V* Benchmark
Val TestA TestB Dev Test Dev Test AR SRR Overall
LLaVA-1.5 + CLIP 61.9 35.7 58.2 76.2 83.4 67.9 65.1 66.5 58.2 58.4 44.4 52.6 47.6
LLaVA-1.5 + SigLIP 2 62.0 37.0 56.6 79.5 84.4 73.9 66.2 64.8 58.0 57.2 40.9 56.6 47.1
LLaVA-1.5 + Meta CLIP 2 62.4 37.0 59.4 76.7 82.8 69.8 66.5 67.0 59.3 60.3 39.1 54.0 45.0
LLaVA-1.5 + FG-CLIP 2 64.0 38.1 62.0 84.9 89.8 79.5 67.6 66.5 60.5 61.4 45.2 57.9 50.3
Table 8: Ablation study results for the training objectives of FG-CLIP 2.
Method COCO80 FG-OVD Flickr30k ImageNet-V2
[…]
FG-CLIP 2 74.9 95.7 52.3 76.3 80.3 92.0 94.1 81.9 72.2
w/o $\mathcal{L}_{\text{CMR}}$ 74.0 94.9 50.9 75.1 77.8 93.5 93.3 81.9 72.1
w/o $\mathcal{L}_{\text{TIC}}$ 70.1 94.8 51.6 75.1 79.1 92.1 93.7 81.8 72.1
w/o $\mathcal{L}_{\text{CMR}}$, $\mathcal{L}_{\text{TIC}}$ 62.7 93.4 51.7 76.0 77.6 90.7 93.5 81.6 72.0
4.3.2 Zero-shot Image Classification
We evaluate zero-shot image classification on ImageNet-1K (Deng et al., 2009) and ImageNet-v2 (Recht et al., 2019) using standard prompts (Radford et al., 2021). As shown in Table 4, FG-CLIP 2 achieves competitive performance compared to SigLIP 2 (Tschannen et al., 2025), and outperforms EVA-CLIP (Sun et al., 2023), Long-CLIP (Zhang et al., 2024a), FineCLIP (Jing et al., 2024), UMG-CLIP (Shi et al., 2024), CLOC (Chen et al., 2024a), and Meta CLIP 2. This confirms that the improvements in fine-grained alignment do not come at the cost of standard recognition accuracy, demonstrating a well-balanced representation capability.
@@ 4.5 Large Multimodal Model tasks
We empoly FG-CLIP 2 as a vision encoder in large multimodal models (LMMs) to assess its compatibility and utility in advanced multimodal reasoning. We integrate FG-CLIP 2 into a standard LLaVA-style LMM architecture, following the pre-training and supervised fine-tuning protocol of LLaVA-1.5 (Liu et al., 2023a). Evaluation is conducted on a diverse set of benchmarks spanning visual question answering (GQA (Hudson & Manning, 2019), TextVQA (Singh et al., 2019)), multimodal understanding (MMMU (Yue et al., 2024)), general multimodal evaluation (MMBench-EN and MMBench-CN (Liu et al., 2024)), referring expression comprehension (RefCOCO (Kazemzadeh et al., 2014)), and fine-grained visual reasoning (V* Benchmark (Wu & Xie, 2024)). Results in Table 7 show that LMMs equipped with FG-CLIP 2 outperform those using other open-source encoders, including SigLIP 2 and Meta CLIP 2. This indicates that the fine-grained and bilingual capabilities of FG-CLIP 2 effectively transfer to higher-level multimodal tasks, making it a strong candidate for integration into next-generation LMMs.
@@ 4.6 Ablation Study
As our training framework leverages global-level, region-level, and hard-negative data through complementary learning mechanisms, we adopt a baseline model that incorporates global alignment alongside fine-grained visual and textual learning. This allows us to specifically evaluate the impact of the Cross-modal Rank Loss with Global Threshold Synchronization ($\mathcal{L}_{\text{CMR}}$) and the Textual Intra-modal Contrastive Loss ($\mathcal{L}_{\text{TIC}}$). As shown in Table 8, removing $\mathcal{L}_{\text{TIC}}$ leads to a 4.8-point drop in COCO Top-1 accuracy (to 70.1%) and a decrease in FG-OVD Hard performance from 52.3% to 51.6%, confirming its critical role in distinguishing semantically similar texts. Removing $\mathcal{L}_{\text{CMR}}$ also degrades performance, with FG-OVD Hard falling to 50.9%, indicating its importance in cross-modal alignment. When both losses are removed, COCO Top-1 drops sharply to 62.7%, demonstrating their complementary benefits. The full model achieves consistent gains across all tasks, especially in bounding box classification and fine-grained understanding, while maintaining strong results on standard recognition benchmarks, validating the effectiveness of our proposed training objectives. Additionally, visualizations demonstrating the semantic separability improvement from TIC loss are provided in Appendix H.
To examine how different data compositions affect the model’s bilingual capability, we conduct an ablation study comparing two variants trained under identical settings except for the data used in the second stage: one uses English-only data, while the other uses both English and Chinese data. Results in Table C (Appendix D) show that the bilingual variant not only achieves higher accuracy on Chinese benchmarks but also consistently improves performance on English-only evaluation sets, confirming a mutually promoting effect between the two languages. We further conduct a comparison with re-trained baselines under identical training data in Appendix E, showing that FG-CLIP 2’s architecture and learning objectives contribute significantly to its performance gains.
@@ 5 Conclusion and Limitations
In this work, we present FG-CLIP 2, a bilingual vision-language model that advances fine-grained understanding for both English and Chinese. Our two-stage training paradigm progressively refines alignment by leveraging both short and long captions, region-text supervision, and multiple discriminative objectives, including the proposed Textual Intra-modal Contrastive (TIC) loss to better distinguish semantically similar descriptions. Trained on large-scale, high-quality English and Chinese datasets, FG-CLIP 2 achieves superior performance across 29 datasets and 8 tasks, demonstrating strong bilingual generalization. To advance evaluation in non-English settings, we introduce a new benchmark for Chinese multimodal understanding with challenging tasks such as long caption image-text retrieval and bounding box classification. We release the model, code, and benchmark to support future research in bilingual fine-grained vision-language understanding. In future work, we will extend the model to handle longer textual inputs and explicitly model relational structures among objects.
Impact Statement
This paper aims to advance the field of Machine Learning, which has broad implications for society. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Training Data Details
Table A summarizes the datasets used in our two-stage training pipeline. We present examples of the Chinese and English training data used in Stage II in Figure A.
@@
FineRegion-CN Chinese 12M
Figure A: Examples of Bilingual Training Data.
@@
Appendix B Open-Vocabulary Object Detection Experimental Details
For each detected bounding box, we extract its visual feature by applying RoI-Align on the dense ViT feature map produced by FG-CLIP 2. We also encode all candidate category names into text embeddings using the same model. We then compute the cosine similarity between the region’s visual feature and each category’s text embedding, followed by Softmax normalization to obtain a category-wise alignment similarity distribution. To produce the final prediction, we combine LLMDet’s original confidence score with FG-CLIP 2’s normalized similarity score via geometric averaging. The resulting fused score reflects both the detector’s localization confidence and the alignment model’s semantic relevance. We then select the category with the highest fused score as the final predicted class, and assign the corresponding fused value as the final confidence output.
@@ Appendix C Dense Prediction Tasks Experimental Details
Appendix D Ablation on Bilingual Data Composition
We present the results of our ablation study on bilingual data composition in Stage II. Both models are initialized from the same Stage I checkpoint trained on the full bilingual (English and Chinese) dataset. In Stage II, we compare two settings: (1) English-only training data, and (2) bilingual English-Chinese training data. As shown in Table C, the bilingual variant not only outperforms the English-only variant on Chinese benchmarks but also consistently improves performance on English-only evaluation sets, confirming a mutually promoting effect between the two languages.
Table C: Ablation study on data composition for bilingual capability. FG-CLIP 2 (marked with ) uses English-only data in Stage II, while FG-CLIP 2 uses both English and Chinese data.
Method Backbone DCI MSCOCO DOCCI-CN Flickr30k-CNA COCO80 Fine-Grained Understanding
[…]
FG-CLIP 2 ViT-B/16 64.4 / 64.7 71.4 / 54.5 57.8 / 58.6 84.5 / 68.0 71.3 / 95.2 52.2 / 75.9 / 80.1 / 91.0
FG-CLIP 2 ViT-B/16 64.5 / 64.9 72.1 / 54.5 71.2 / 75.4 85.4 / 69.9 74.9 / 95.7 52.3 / 76.3 / 80.3 / 92.0
Appendix E Comparison with Re-trained Baselines
To ensure a fair comparison and isolate the impact of model architecture and learning objectives from that of training data, we re-train two representative baselines, SigLIP 2 and Chinese-CLIP, using exactly the same Stage II training data as FG-CLIP 2. Due to computational constraints, all models in this study, including FG-CLIP 2, are trained solely on the Stage II dataset without any Stage I pretraining. This setup guarantees identical data conditions across all methods, enabling a fully controlled evaluation focused solely on architectural and objective-level differences.
Table D summarizes the results under this setting. Remarkably, even in the absence of large-scale pretraining, FG-CLIP 2 consistently outperforms both SigLIP 2 and Chinese-CLIP across all benchmarks, including DCI, MSCOCO, DOCCI-CN, Flickr30k-CNA, COCO80, and Fine-Grained Understanding. These results indicate that, beyond the benefits of bilingual training data, the architectural design and tailored learning objectives of FG-CLIP 2 play a crucial role in its superior performance. The consistent gains across diverse benchmarks demonstrate that these modeling choices substantially enhance the model’s ability to capture fine-grained semantics and align cross-modal representations effectively.
Table D: Comparison with retrained baselines using only Stage II data (no Stage I pretraining, indicated by †).
Method Backbone DCI MSCOCO DOCCI-CN Flickr30k-CNA COCO80 Fine-Grained Understanding
[…]
Chinese-CLIP† ViT-B/16 26.9 / 25.5 47.3 / 31.7 48.3 / 47.7 80.0 / 65.5 41.4 / 69.0 16.3 / 33.0 / 37.1 / 73.9
SigLIP 2† ViT-B/16 49.1 / 49.0 71.5 / 53.5 50.5 / 49.4 81.0 / 52.7 53.7 / 79.7 25.5 / 50.6 / 59.2 / 83.4
FG-CLIP 2† ViT-B/16 60.9 / 62.6 72.1 / 53.7 65.9 / 69.5 84.3 / 66.1 71.0 / 95.4 51.9 / 76.0 / 80.5 / 90.6
Appendix F Hyperparameter Ablation Study
In addition to the global alignment learning objective, our model incorporates four auxiliary losses: Fine-Grained Visual Learning (FGV), Fine-Grained Textual Learning (FGT), Textual Intra-modal Contrastive Loss (TIC), and Cross-Modal Rank Loss (CMR). The weights for these losses are set using a sequential tuning strategy: starting from the base model trained with only the global alignment loss, we introduce each auxiliary loss one at a time, tune its weight while keeping previously added losses fixed (and excluding any future ones), and select the value that yields the highest average performance on the validation set.
Figure B shows the average performance over different values of each loss hyperparameter, where the average is computed across several representative evaluation tasks: DCI, MSCOCO, DOCCI-CN, Flickr30k-CNA, Bbox Classification (COCO80), and Fine-Grained Understanding.
Figure B: The performance over different values of each loss hyperparameter.